OpenStudy (anonymous):

Randy wants to move his savings account to a new bank that pays a better interest rate of 3.5 percent so that he can earn 100 dollars in interest faster that at his old bank. If he moves 800 dollars to the new bank, how long will it take for him to earn the 100 dollars in interest?

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

3.5 % of $800 = 0.035 *800 =28 $ now number of days = 100/28

OpenStudy (anonymous):

tnx! so i just divide?

OpenStudy (ghazi):

of course :)

OpenStudy (anonymous):

got it:)

OpenStudy (anonymous):

tnx again!

OpenStudy (ghazi):

:D YW

OpenStudy (anonymous):

another one coming up!

OpenStudy (anonymous):

(a question)

OpenStudy (ghazi):

alright its okay :D

OpenStudy (anonymous):

gur eight!

OpenStudy (ghazi):

is that your question?

OpenStudy (anonymous):

nope. 'gr8' lol

OpenStudy (ghazi):

oh god :)